# Mind-to-Image: Projecting Visual Mental Imagination of the Brain from fMRI

### v4

@@ Abstract
Machine Learning, ICML
[…]
@@ 1 Introduction
Figure 1: Overview of our Mind-to-Image pipeline which allows to reconstruct images from mental imagery. Left: overview of MindEye, the fMRI-to-Image model used in our Mind-to-Image approach. The stimuli (an image seen or in our case imagined) and the associated fMRI data is fed to a high-level pipeline (composed of an MLP and Diffusion Prior) and a low-level pipeline to create aligned CLIP embeddings of the fMRI data. Middle: fMRI data collection protocols. In the classic fMRI-to-Image approach, a subject sees images in an MRI scanner, BOLD (Blood-Oxygen Level Dependent) data are recorded then used to train an fMRI-to-Image model that convert brain data to match the seen images. In our approach, we devise two mental imagery protocols: weak and strong imagination. For weak imagination, we let the subject imagine previously seen images from a dataset of surrealist images (face portraits and nature landscapes) which we create. The brain data associated to the recollection of such images is gathered. For strong imagination, we collect brain data where the subject completely imagines new images based on instructions. Right: our Mind-to-Image pipeline. We use the weak imagination fMRI data to train an fMRI-to-Image model to reconstruct recollected images from visual imagination brain data. At inference time we use strong imagination fMRI data and this trained model to generate the imagined images, thanks to transfer learning.
@@
In the rapidly evolving field of neuroscience and artificial intelligence, the quest to decode the complex mechanisms of the human brain has led to groundbreaking developments, particularly in the area of functional magnetic resonance imaging (fMRI). Among the myriad applications of fMRI, one of the most intriguing has been the reconstruction of visual experiences. Traditional methods have successfully reconstructed images based on fMRI BOLD (blood-oxygen-level-dependent) data captured while subjects view specific images (Chen et al., 2022, 2023; Scotti et al., 2023). This process involves preprocessing the BOLD data to isolate voxels associated with the brain’s visual areas, generating clip embeddings from these voxels, and then employing image generation algorithms that utilize these embeddings to recreate the observed images. Such methodologies have shown promising results, especially when employing datasets like the NSD (Natural Scenes Dataset) (Allen et al., 2022), a substantial 7T fMRI dataset designed to bridge cognitive neuroscience and artificial intelligence. This dataset, featuring extensive scans from eight subjects, has had tremendous impact for research in this domain.
@@
Progress in cognitive neuroscience and artificial intelligence has not only allowed to advance our understanding of brain mechanisms but also to holds the promise to develop practical applications, such as real-time clinical tools and brain-computer interfaces. However, despite the progress in visual reconstruction from observed images, a challenge remains in the domain of visual imagination. Visual imagination encompasses both the recall of previously seen images (which we refer to as weak imagination in the paper) and the generation of entirely new images within the mind’s eye (which we refer to as strong imagination). The ability to decode and reconstruct these imagined visuals offers a wide range of promising possibilities from aiding individuals with disabilities, to enhancing creative processes.
@@
Leveraging the MindEye model (Scotti et al., 2023), a state-of-the-art fMRI-to-image framework, we adapt its architecture to adapt to the larger dimensional complexity inherent in visual imagination data. Our approach is summarised in Fig.1. Our results, while preliminary, indicate the model’s capability to, at least to some degree, generate portraits and landscapes based on the nature of the imagination. Although the fidelity of the reconstructions to the imagined content varies, which is measured quantitatively through usual metrics in the field, our findings underscore the potential of this approach. With further refinement and expansion of the dataset, this research could pave the way for more precise reconstructions of visual imagination, with the promise of a better understanding of the brain’s creative and cognitive processes.
@@ 2 Related work
In the domain of mental imagery, several studies have sought to decode and reconstruct images from the brain’s visual cortex when subjects engage in imagining specific scenes or objects. (Stokes et al., 2009) provided early evidence of the brain’s capacity to activate shape-specific population codes during mental imagery, highlighting the top-down influence of visual imagination on neural activity in the visual cortex. (Reddy et al., 2010) explored the decoding of category information during mental imagery, providing an understanding on how different types of visual information are processed and represented in the brain.
More recently, (Naselaris et al., 2015) developed a voxel-wise encoding model capable of decoding mental images of remembered scenes, demonstrating the relationship between memory, imagination, and visual perception. (VanRullen & Reddy, 2019) employs a combination of variational auto-encoders (VAE) and Generative Adversarial Networks (GAN) to perform pairwise decoding of faces with a high degree of accuracy but also to classify gender and decode which face was imagined by subjects. (Goebel et al., 2022) showed the feasibility of reading imagined letter shapes from the mind’s eye using real-time 7 Tesla fMRI, presenting a novel approach to understanding how specific visual information can be decoded from brain activity.
@@
More close to our work, a recent paper (Koide-Majima et al., 2024) provides preliminary results on reconstructing images from memory. While their results are encouraging, their reconstruction pipeline backbone is not state-of-the-art like MindEye is, which partly explains the results obtained.
@@
The development of a comprehensive data collection protocol is one the main contributions of our study, aimed at capturing the complex neural correlates of visual imagination. Drawing inspiration from the NSD paper (Allen et al., 2022), our protocol is designed to explore two distinct modes of visual imagination: weak imagination (imagination from memory) and strong imagination (pure imagination). This approach enables us to study how the brain navigates between recalling visual information and generating novel visual constructs.
@@ 3.1 Image dataset used: Surrealism
We do not use the images from the COCO dataset (Lin et al., 2014) as it is done in the NSD dataset. Because our research is also associated to an artistic research, we decided to narrow the dataset to one theme (surrealism) and two modes (face portraits and nature landscapes). This choice narrows the visual space of the fMRI-to-Image model, which simplify the task and allows us to have results without having to do 40 hours of scans as in the NSD dataset.
Thus, we created our own dataset of images for our experiments. The images come from three sources: real images and generated images using Versatile Diffusion (Xu et al., 2023) and Midjourney. The dataset comprises 1200 images.
@@
Data were collected on Siemens 3T PrismaFit using a 64 channels headcoil. BOLD data were acquired using the CMRR multiband EPI sequence. The main parameters are: 2mm isotropic voxels with full brain coverage, TR/TE = 1300ms/27ms, MultiBand=4, partialFourier=7/8, FlipAngle=68°, P>>much-greater-than>>> >A phase encoding direction, EchoSpacing at minimum.
@@ 3.4 Strong imagination: data collection protocol
For strong imagination, the subject is now asked to generate entirely new visual content based on verbal prompts. The subject is instructed to imagine a portrait or landscape evoking one of ten primal emotions (such as fear, happiness, love, etc.). This task is designed to stimulate the pure imagination process, encouraging the brain to construct novel visual imagery without the reference of previously seen images. Each imagination session lasts for 6 seconds, followed by a 1-second rest, cycling through combinations of landscape/portrait and emotional prompts. This cycle is repeated 10 times. Unlike the data collected for weak imagination, this dataset is solely used during the model’s inference stage. The evaluation here is similar to the weak imagination evaluation: quantitively assess if the model matches the correct modality (face portrait or nature landscape) and qualitatively assess if the contents of the generated image matches the visual imagery of the subject. However we cannot compute all the quantitative metrics as the ground truth image does not actually exists outside the subject’s mind. To ensure that we recall what the subject has imagined, the subject provides an oral description of the visual contents of the imagined image.
@@ 3.5 Pre-processing of raw BOLD data and masking of the regions of interest
After the raw BOLD (blood-oxygen-level-dependent) data is collected, we apply a preprocessing routine to provide the model with data that decorrelates all the noise that impair the following steps. We first co-register all BOLD data with a T1 anatomical scan. Then, using a General Linear Model (GLM) with the Glover hemodynamic response function (HRF) (implemented via the Nilearn library (Abraham et al., 2014)), we extracted beta values associated with each voxel for every distinct event, including both the displayed images and the periods of imagination. This allows us to isolate the specific brain activity corresponding to each event, which is common practice for such tasks.
In the subsequent data pre-processing phase, we identify the brain regions most important to visual imagination and encoding, using a GLM model to manually extract these areas of interest. Unlike the NSD dataset, which predominantly focuses on visual areas, our study required the creation of a custom mask to englobe the regions identified as crucial in our context of weak and strong imagination. This aforementioned mask was manually curated based on the GLM analysis results, ensuring that only voxels corresponding to our defined regions of interest were included in further analyses. The delineation and analysis of these regions, and how they diverge from the predominantly visual areas highlighted in the NSD study, is studied in the results section.
Finally, we create our training, validation (weak imagination) and test (strong imagination) datasets by flattening the pre-processed betas corresponding the selected voxels and pair them with the images or prompts related to their events.
@@ 4 Mind-to-Image pipeline
We know describe the fMRI-to-Image model used for our experiments and then outline our Mind-to-Image pipeline by connecting the dots between the data collection protocol and the fMRI-to-Image model.
@@
We start from the state-of-the-art MindEye model which reconstructs images viewed by subjects from fMRI data by combining two pipelines: a high-level (semantic) pipeline and a low-level (perceptual) pipeline. The high-level pipeline maps fMRI voxels to the CLIP Vision Transformer L/14 image space, capturing the semantics of the image. The low-level pipeline targets the perceptual fidelity of reconstructed images by mapping voxels to Stable Diffusion’s Variational Autoencoder (VAE) embedding space. They leverage contrastive and mean squared error (MSE) losses for training of its projection and reconstruction capabilities. At inference time, fMRI data associated to an image is fed through the model which outputs a low-level reconstructed image and aligned CLIP embeddings. These embeddings are fed to an UnCLIP models, which is a model able to reconstruct an image based on its CLIP embeddings. In this case, the embeddings as well as the low-level reconstructed image is fed to a Versatile Diffusion Image Variations pipeline to obtain the final reconstruction, as it is done in the original MindEye paper. The whole fMRI-to-Image pipeline is illustrated in Fig.1.
@@ 5 Results
We first analyze the most activate regions of the brain during visual perception, weak imagination and strong imagination.
@@ 5.1 What happens in the brain during visualisation, weak and strong imagination?
In this section we describe the brain regions involved during visual perception, weak imagination, and strong imagination. This is done though the visualisations presented in Fig.2.
Figure 2: Functional MRI results of the three conditions, respectively: simple visual perception (A- in red), weak imagination (B- in pink) and strong imagination (C- in blue) presented through a lower (top row) and right lateral (bottom row) view. Activation maps were generated by comparing the “test” and “rest” conditions, using the same statistical threshold for each test (T-score =5).
@@ 5.2 Can we reconstruct images from weak imagination / memory?
Figure 3: Results for weak imagination on the validation dataset. Top row: images from the validation set of our surrealism image dataset, which were recollected from memory by the subject in the fMRI scanner. Bottom row: reconstructed images from the fMRI-to-Image model based on the brain data associated to the recollection of the images in the top row.
Method Low-Level High-Level
PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
@@
Table 1: Quantitative comparison of Mind-to-Image reconstruction performance against other models trained on different datasets, computed on usual metrics in the field. The results are not directly comparable, but are given to provide an element of comparison on the ranges of the values of the metrics. PixCorr is the pixel-level correlation of reconstructed and groundtruth images. SSIM (Wang et al., 2004) is the structural similarity index metric. AlexNet(2) and AlexNet(5) are the 2-way comparisons of the second and fifth layers of AlexNet (Krizhevsky et al., 2012), respectively. Inception is the 2-way comparison of the last pooling layer of InceptionV3 (Szegedy et al., 2016). CLIP is the 2-way comparison of the output layer of the CLIP-Vision model (Radford et al., 2021). EffNet-B and SwAV are distance metrics gathered from EfficientNet-B (Tan & Le, 2019) and SwAV-ResNet (Caron et al., 2020) models, respectively.
In the exploration of weak imagination, where the subject is asked to recall previously seen images, our results show the capacity of the model to identify to the original class of the images. Specifically, if a subject recalled a face portrait, the reconstructed image is almost always a face portrait, and similarly for nature landscapes. This observation suggests a the ability of the model to grasp the category of the imagined content accurately within the BOLD data.
First, to quantitatively assess this capability, we fine-tune a pretrained ResNet50 model (He et al., 2016) to classify images as either face portraits or nature landscapes using the images in our dataset. We apply this classifier to the reconstructed images from the weak imagination tasks of the validation set. We obtain a classification accuracy of 91% in correctly identifying the reconstructed images as belonging to their original categories of either portraits or landscapes. This level of accuracy demonstrates the model’s ability in capturing the category information within the neural activity during weak imagination.
@@
Then, we provide in Tab.1 the usual metrics to quantitatively assess the performance of reconstruction from weak imagination. We compare our results with MindEye (Scotti et al., 2023) Note that the comparisons are not apples-to-apples: (Scotti et al., 2023) provide results for visual stimuli experiments (not for weak imagination) on a large-scale (approx. 10x size) dataset that uses another image domain (realistic images), explaining the big gap in performance. Our results, while inferior to those of reconstructions from visual stimuli, show that we are able to have significantly better than chance results on reconstruction from weak imagination. We noted two main outliers on our quantitative results: we get poor results on the SSIM metric, and excellent results on the SwAV metric. We posit that the difference comes from the nature and diversity of images we used (narrow surrealist images compared to more various natural images), but further investigation remain needed to better understand this surprising result.
@@
Qualitatively, while the model demonstrated a robust ability to regenerate the broad category of the original images, the fidelity of content reconstruction within these images presented a more varied outcome. Some reconstructions were visually semantically close to the remembered images. However, others diverged more significantly, indicating a variability in the model’s ability to precisely capture and reconstruct the nuanced details of each imagined scene. We provide uncurated reconstructions in Fig.5 in Appendix. This variance in reconstruction fidelity can be attributed, in part, to the limitations imposed by the dataset’s size and the duration of the fMRI scanning sessions. With only 6 hours of scans and 1200 images constituting the dataset, the goal for training a model from scratch to achieve perfect reconstruction is out of reach. Nevertheless, these initial results are promising, suggesting a capacity for accurate content reproduction. It is reasonable to anticipate that with an expanded dataset, the fidelity of the reconstructed images to their original counterparts would see important improvement, enhancing both the precision and consistency of the content reproduction in future iterations.

### v5

@@ Abstract
color=red!60,inline]H:TODO list:Better figures and explanations for charles part, other approach: baseline with mindeye-v2 fine tuning, appendix: hyperparams
@@ 1 Introduction
Fig. 1: Overview of our Mind-to-Image pipeline which allows to reconstruct images from mental imagery. Left: overview of MindEye, the fMRI-to-Image model used in our Mind-to-Image approach. The stimuli (an image seen or in our case imagined) and the associated fMRI data is fed to a high-level pipeline (composed of an MLP and Diffusion Prior) and a low-level pipeline to create aligned CLIP embeddings of the fMRI data. Middle: fMRI data collection protocols. In the classic fMRI-to-Image approach, a subject sees images in an MRI scanner, BOLD (Blood-Oxygen Level Dependent) data are recorded then used to train an fMRI-to-Image model that converts brain data to match the seen images. In our approach, we devise two mental imagery protocols: weak and strong imagination. For weak imagination, we let the subject imagine previously seen images from a dataset of surrealist images (face portraits and nature landscapes) which we create. The brain data associated with the recollection of such images is gathered. For strong imagination, we collect brain data where the subject completely imagines new images based on instructions. Right: our Mind-to-Image pipeline. We use the weak imagination fMRI data to train an fMRI-to-Image model to reconstruct recollected images from visual imagination brain data. At inference time we use strong imagination fMRI data and this trained model to generate the imagined images, thanks to transfer learning.
@@
In the rapidly evolving field of neuroscience and artificial intelligence, the quest to decode the complex mechanisms of the human brain has led to groundbreaking developments, particularly in the area of functional magnetic resonance imaging (fMRI). Among the myriad applications of fMRI, one of the most intriguing has been the reconstruction of visual experiences. Traditional methods have successfully reconstructed images based on fMRI BOLD (blood-oxygen-level-dependent) data captured while subjects view specific images [5, 6, 16]. This process involves preprocessing the BOLD data to isolate voxels associated with the brain’s visual areas, generating clip embeddings from these voxels, and then employing image generation algorithms that utilize these embeddings to recreate the observed images. Such methodologies have shown promising results, especially when employing datasets like the NSD (Natural Scenes Dataset) [2], a substantial 7T fMRI dataset designed to bridge cognitive neuroscience and artificial intelligence. This dataset, featuring extensive scans from eight subjects, has had tremendous impact for research in this domain.
@@
Progress in cognitive neuroscience and artificial intelligence has not only allowed to advance our understanding of brain mechanisms but also to hold the promise to develop practical applications, such as real-time clinical tools and brain-computer interfaces. However, despite the progress in visual reconstruction from observed images, a challenge remains in the domain of visual imagination. Visual imagination encompasses both the recall of previously seen images (which we refer to as weak imagination in the paper) and the generation of entirely new images within the mind’s eye (which we refer to as strong imagination). The ability to decode and reconstruct these imagined visuals offers a wide range of promising possibilities from aiding individuals with disabilities, to enhancing creative processes.
@@
Leveraging the MindEye model [16], a state-of-the-art fMRI-to-image framework, we adapt its architecture to adapt to the larger dimensional complexity inherent in visual imagination data. Our approach is summarised in Fig.1. Our results, while preliminary, indicate the model’s capability to, at least to some degree, generate portraits and landscapes based on the nature of the imagination. Although the fidelity of the reconstructions to the imagined content varies, which is measured quantitatively through usual metrics in the field, our findings underscore the potential of this approach. With further refinement and expansion of the dataset, this research could pave the way for more precise reconstructions of visual imagination, with the promise of a better understanding of the brain’s creative and cognitive processes.
@@ 2 Related work
In the domain of mental imagery, several studies have sought to decode and reconstruct images from the brain’s visual cortex when subjects engage in imagining specific scenes or objects. [18] provided early evidence of the brain’s capacity to activate shape-specific population codes during mental imagery, highlighting the top-down influence of visual imagination on neural activity in the visual cortex. [15] explored the decoding of category information during mental imagery, providing an understanding on how different types of visual information are processed and represented in the brain.
More recently, [13] developed a voxel-wise encoding model capable of decoding mental images of remembered scenes, demonstrating the relationship between memory, imagination, and visual perception. [21] employs a combination of variational auto-encoders (VAE) and Generative Adversarial Networks (GAN) to perform pairwise decoding of faces with a high degree of accuracy but also to classify gender and decode which face was imagined by subjects. [8] showed the feasibility of reading imagined letter shapes from the mind’s eye using real-time 7 Tesla fMRI, presenting a novel approach to understanding how specific visual information can be decoded from brain activity.
@@
More close to our work, a recent paper [10] provides preliminary results on reconstructing images from memory. While their results are encouraging, their reconstruction pipeline backbone is not state-of-the-art like MindEye is, which partly explains the results obtained.
@@
The development of a comprehensive data collection protocol is one the main contributions of our study, aimed at capturing the complex neural correlates of visual imagination. Drawing inspiration from the NSD paper [2], our protocol is designed to explore two distinct modes of visual imagination: weak imagination (imagination from memory) and strong imagination (pure imagination). This approach enables us to study how the brain navigates between recalling visual information and generating novel visual constructs.
@@ 3.1 Image dataset used: Surrealism
We do not use the images from the COCO dataset [12] as it is done in the NSD dataset. Because our research is also associated with artistic research, we decided to narrow the dataset to one theme (surrealism) and two modes (face portraits and nature landscapes). This choice narrows the visual space of the fMRI-to-Image model, which simplifies the task and allows us to have results without having to do 40 hours of scans as in the NSD dataset.
Thus, we created our own dataset of images for our experiments. The images come from three sources: real images and generated images using Versatile Diffusion [23] and Midjourney. The dataset comprises 1200 images.
@@
Data was collected on Siemens 3T PrismaFit using a 64 channels headcoil. BOLD data were acquired using the CMRR multiband EPI sequence. The main parameters are: 2mm isotropic voxels with full brain coverage, TR/TE = 1300ms/27ms, MultiBand=4, partialFourier=7/8, FlipAngle=68°, P>>much-greater-than$>$ >A phase encoding direction, EchoSpacing at minimum.
@@ 3.4 Strong imagination: data collection protocol
For strong imagination, the subject is now asked to generate entirely new visual content based on verbal prompts. The subject is instructed to imagine a portrait or landscape evoking one of ten primal emotions (such as fear, happiness, love, etc.). This task is designed to stimulate the pure imagination process, encouraging the brain to construct novel visual imagery without the reference of previously seen images. Each imagination session lasts for 6 seconds, followed by a 1-second rest, cycling through combinations of landscape/portrait and emotional prompts. This cycle is repeated 10 times. Unlike the data collected for weak imagination, this dataset is solely used during the model’s inference stage. The evaluation here is similar to the weak imagination evaluation: quantitatively assess if the model matches the correct modality (face portrait or nature landscape) and qualitatively assess if the contents of the generated image matches the visual imagery of the subject. However we cannot compute all the quantitative metrics as the ground truth image does not actually exist outside the subject’s mind. To ensure that we recall what the subject has imagined, the subject provides an oral description of the visual contents of the imagined image.
@@ 3.5 Pre-processing of raw BOLD data and masking of the regions of interest
After the raw BOLD (blood-oxygen-level-dependent) data is collected, we apply a preprocessing routine to provide the model with data that decorrelates all the noise that impair the following steps. We first co-register all BOLD data with a T1 anatomical scan. Then, using a General Linear Model (GLM) with the Glover hemodynamic response function (HRF) (implemented via the Nilearn library [1]), we extracted beta values associated with each voxel for every distinct event, including both the displayed images and the periods of imagination. This allows us to isolate the specific brain activity corresponding to each event, which is common practice for such tasks.
Mask selection. In the subsequent data pre-processing phase, we identify the brain regions most important to visual imagination and visual perception, using a GLM model to manually extract these areas of interest. Unlike the NSD dataset, which predominantly focuses on visual areas, our study required the creation of a custom mask to englobe the regions identified as crucial in our context of weak and strong imagination.
It is a well known fact that the generation of the perceptive or imagined visual experience [7] involves the ventral stream cortical areas (in occipital and temporal lobe). In these visual regions, various aspects of visual information contribute to creating a mental image. This includes features like shape, color, motion, depth, and texture. These features are processed and integrated across low-level (V1,V2), associative (V3, V4, V5) and high-level (including fusiform gyrus in temporal lobe) cortical areas to form a coherent mental representation of what we see or what we think of.
This is the reason why we designed the brain mask to collect the cortical responses during the imagination conditions using the mean responses of the general linear model applied to the event vs. rest condition, in the occipital and temporal lobe. This method aimed to consider only the cortical activity involved in creating a mental image in one subject during the imagination phase.
We thus create and compare three masks manually curated based on the GLM analysis results, ensuring that only voxels corresponding to our defined regions of interest were included in further analyses. The delineation and analysis of these regions, and how they diverge from the predominantly visual areas highlighted in the NSD study, is studied in the results section.
Finally, we create our training, validation (weak imagination) and test (strong imagination) datasets by flattening the pre-processed betas corresponding to the selected voxels and pair them with the images or prompts related to their events.
@@ 4 Mind-to-Image pipeline
We now describe the fMRI-to-Image model used for our experiments and then outline our Mind-to-Image pipeline by connecting the dots between the data collection protocol and the fMRI-to-Image model.
@@
We experiment with two approaches for the choice of the fMRI-to-Image model using in our Mind-to-Image pipeline.
4.1.1 Specialized model approach
We start from the state-of-the-art MindEye model which reconstructs images viewed by subjects from fMRI data by combining two pipelines: a high-level (semantic) pipeline and a low-level (perceptual) pipeline. The high-level pipeline maps fMRI voxels to the CLIP Vision Transformer L/14 image space, capturing the semantics of the image. The low-level pipeline targets the perceptual fidelity of reconstructed images by mapping voxels to Stable Diffusion’s Variational Autoencoder (VAE) embedding space. They leverage contrastive and mean squared error (MSE) losses for training of its projection and reconstruction capabilities. At inference time, fMRI data associated to an image is fed through the model which outputs a low-level reconstructed image and aligned CLIP embeddings. These embeddings are fed to an UnCLIP model, which is a model able to reconstruct an image based on its CLIP embeddings. In this case, the embeddings as well as the low-level reconstructed image is fed to a Versatile Diffusion Image Variations pipeline to obtain the final reconstruction, as it is done in the original MindEye paper. The whole fMRI-to-Image pipeline is illustrated in Fig.1.
@@ 4.1.1 Specialized model approach
4.1.2 Alternative approach: fine-tuning from pre-trained fMRI-to-Image model on vision.
Another approach we experimented with is to leverage pre-trained multi-subject models on a vision task on the NSD dataset. Then, we hope that fine tuning these models with our imagination data will help the model generalize better, motivated by the overlap between the vision and imagination tasks. We thus experiment with MindEye2 [17], the recent update to MindEye which among several technical improvements provides a multi-subject model on the NSD dataset (vision task). This model can be fine-tuned given user data, which we proceed to do in our experiments. However there are a few issues to consider: the data is not collected the same way as the NSD data, nor pre-processed exactly the same, which we posit can hinder the approach.
@@ 5 Results
We first analyze the most activated regions of the brain during visual perception, weak imagination and strong imagination.
@@ 5.1 What happens in the brain during visualisation, weak and strong imagination?
In this section we describe the brain regions involved during visual perception, weak imagination, and strong imagination. This is done through the visualisations presented in Fig.2.
Fig. 2: Functional MRI results of the three conditions, respectively: simple visual perception (A- in red), weak imagination (B- in pink) and strong imagination (C- in blue) presented through a lower (top row) and right lateral (bottom row) view. Activation maps were generated by comparing the “test” and “rest” conditions, using the same statistical threshold for each test (T-score =5).
@@ 5.2 Can we reconstruct images from weak imagination / memory?
Stimuli & Dataset Low-Level High-Level
Model PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
Vision results on NSD dataset
MindEye [16] .${.309}.$323.323{.323}.323 ${94.7\%}$ ${97.8\%}$ ${93.8\%}$ ${94.1\%}$${.645}.$367.367{.367}.367
Mind-to-Image
(Weak imagination on Surrealism dataset)
MindEye1 - Mask Vision .${.177}.$051.051{.051}.051 ${57.8\%}$ ${53.0\%}$ ${58.4\%}$ ${57.9\%}$${.718}.$079.079{.079}.079
MindEye1 - Mask Imagination .${.082}.$023.023{.023}.023 ${54.8\%}$ ${51.6\%}$ ${55.1\%}$ ${54.9\%}$${.797}.$098.098{.098}.098
@@
Table 1: Quantitative comparison of Mind-to-Image reconstruction performance against other models trained on different datasets, computed on usual metrics in the field. The results are not directly comparable, but are given to provide an element of comparison on the ranges of the values of the metrics. PixCorr is the pixel-level correlation of reconstructed and groundtruth images. SSIM [22] is the structural similarity index metric. AlexNet(2) and AlexNet(5) are the 2-way comparisons of the second and fifth layers of AlexNet [11], respectively. Inception is the 2-way comparison of the last pooling layer of InceptionV3 [19]. CLIP is the 2-way comparison of the output layer of the CLIP-Vision model [14]. EffNet-B and SwAV are distance metrics gathered from EfficientNet-B [20] and SwAV-ResNet [4] models, respectively.
@@
Quantitative results. We provide in Tab.1 the usual metrics to quantitatively assess the performance of reconstruction from weak imagination. As an indication, we compare our results with MindEye [16]. Note that the comparisons are not apples-to-apples: [16] provide results for visual stimuli experiments (not for weak imagination) on a large-scale (approx. 10x size) dataset that uses another image domain (realistic images), explaining the big gap in performance. Our results, while inferior to those of reconstructions from visual stimuli, show that we are able to have significantly better than chance results on reconstruction from weak imagination.
[…]
Regarding the metric scores, correlations/distance between the reconstructed and original images (PixCorr, SSIM, Eff and SwAV) do not seem representative of model performance compared to 2-way comparisons (Alex(2), Alex(5), Incep and CLIP). For instance, MindEye2 multi-subject fine-tuning scores extremely well on correlations/distance metrics, while qualitatively we observed poor results. MindEye2 uses SD-XL for its image generation, which is stronger than Versatile Diffusion used in MindEye1: this could explain the result as the generated images with MindEye2 are naturally better looking. For clarity purposes, we decided not to bold the results for correlation/distance metrics and based our comparisons on the 2-way comparisons metrics. The quality of metrics is a known problem in the field and some authors have shifted toward human-based assessments, which can be hard to reproduce under the same conditions. Further investigation remains needed to assess the quality of the metrics used in the field.
@@
Qualitative results. While the model demonstrated a robust ability to regenerate the broad category of the original images, the fidelity of content reconstruction within these images presented a more varied outcome. Some reconstructions were visually semantically close to the remembered images. However, others diverged more significantly, indicating a variability in the model’s ability to precisely capture and reconstruct the nuanced details of each imagined scene. We provide uncurated reconstructions in Fig.5 in Appendix. This variance in reconstruction fidelity can be attributed, in part, to the limitations imposed by the dataset’s size and the duration of the fMRI scanning sessions. With only 6 hours of scans and 1200 images constituting the dataset, the goal for training a model from scratch to achieve perfect reconstruction is out of reach. Nevertheless, these initial results are promising, suggesting a capacity for accurate content reproduction. It is reasonable to anticipate that with an expanded dataset, the fidelity of the reconstructed images to their original counterparts would see important improvement, enhancing both the precision and consistency of the content reproduction in future iterations.
Fig. 3: Results for weak imagination on the validation dataset. Top row: images from the validation set of our surrealism image dataset, which were recollected from memory by the subject in the fMRI scanner. Bottom row: reconstructed images from the fMRI-to-Image model based on the brain data associated with the recollection of the images in the top row.
Fig. 4: Results for strong imagination. Subject is given an instruction on the fMRI screen, such as "Imagine a portrait representing optimism". Then, the subject purely imagines such an image, giving an oral description. Then, the associated brain data is fed to the fMRI-to-Image model to produce a reconstruction, evaluated compared to the oral description.